OpenStudy (anonymous):

Multiply 4(x^3 - 2x + 4) (Points : 1) 4x3 - 2x + 4 4x3 - 8x + 16 4x3 + 2x + 8

OpenStudy (owlcoffee):

for the solution of this problem, we shall recall the distributive property, since it's a postulate, it does not have a proof and I can easily write it for you: \[a(b+c)=ab+ac\] What it states, is that if a real number or a variable is multiplying another sum of variables or constants, we can transform it into a sum if we multiply all the terms by the multiplying number or variable. so having: \[4(x ^{3}+2x+4)= 4x ^{3}+8x +16\]

OpenStudy (owlcoffee):

oh, sorry, did not see the "-" sign. the result is: \[4x ^{3}-8x+16\]

OpenStudy (anonymous):

Oh! Ok thanks so much I totally understand that.